# Weak-to-Strong Reasoning

### v1

@@ Abstract
When large language models (LLMs) exceed human-level capabilities, it becomes increasingly challenging to provide full-scale and accurate supervisions for these models. Weak-to-strong learning, which leverages a less capable model to unlock the latent abilities of a stronger model, proves valuable in this context. Yet, the efficacy of this approach for complex reasoning tasks is still untested. Furthermore, tackling reasoning tasks under the weak-to-strong setting currently lacks efficient methods to avoid blindly imitating the weak supervisor including its errors. In this paper, we introduce a progressive learning framework that enables the strong model to autonomously refine its training data, without requiring input from either a more advanced model or human-annotated data. This framework begins with supervised fine-tuning on a selective small but high-quality dataset, followed by preference optimization on contrastive samples identified by the strong model itself. Extensive experiments on the GSM8K and MATH datasets demonstrate that our method significantly enhances the reasoning capabilities of Llama2-70b using three separate weak models. This method is further validated in a forward-looking experimental setup, where Llama3-8b-instruct effectively supervises Llama3-70b on the highly challenging OlympicArena dataset. This work paves the way for a more scalable and sophisticated strategy to enhance AI reasoning powers. All relevant code and resources are available in https://github.com/GAIR-NLP/weak-to-strong-reasoning.
@@ 1 Introduction
To achieve the above goal, we introduce a progressive refinement learning framework, guided by the principle that a model can enhance its capabilities more effectively by initially focusing on smaller, more reliable subsets of data, and then iteratively expanding its learning scope, as illustrated in Fig. 2. In the first stage, we hypothesize that it is more advantageous to utilize smaller quantities of data that are likely to be more accurate. We achieve this by combining weak data, generated by the less capable model, with data self-generated by the more advanced model through in-context learning. This blend is then used to selectively curate datasets for subsequent supervised fine-tuning. In the second stage, upon having developed a strong model with improved reasoning capabilities, we utilize its ability to construct contrastive samples for preference optimization (Rafailov et al., 2023; Hong et al., 2024) and enables the model to learn effectively from the errors of the weaker model.
@@
To more accurately approximate future scenarios, we conduct additionally experiments on OlympicArena (Huang et al., 2024), an extremely challenging dataset with no definitive ground truth answers. Llama3-8b-instruct (AI@Meta, 2024), despite its smaller size, has been aligned and proved to effectively supervise the larger Llama3-70b, whose potential have not yet been fully realized. Moreover, our proposed two-stage training approach outperforms full weak fine-tuning by 3.19 points.
@@ G.T. Answer Stronger Model
Gaining insights from semi-supervised learning within the domain of traditional machine learning, another type of LLM learning depends not on extensive labeling but instead on a small, high-quality seed dataset. Tong et al. (2024) have demonstrated improvement by learning differences between self-generated responses and expert-annotated responses. We also include the trending research topic of easy-to-hard generalization (Hase et al., 2024; Sun et al., 2024) in this category, where models are trained to tackle complex tasks by learning from human annotations on easier tasks. This series of research inevitably require access to a small yet high quality set of standard answers.
@@ 2.2 Weak-to-Strong Reasoning Setup
In this paper, we address reasoning tasks in the weak-to-strong setting, as illustrated in Tab. 2. First, we examine mathematical reasoning tasks, such as those in GSM8k and MATH. These tasks require each step of the reasoning process to demonstrate fundamental mathematical problem-solving skills, including problem comprehension and algebraic operations, and build upon the previous steps. It imposes higher demands on the model’s learning and generalization capabilities. Unlike classification tasks, where models can rely on superficial pattern extrapolation or recognition, reasoning tasks offer minimal benefit from guessing. Then, we use a weak model (e.g., Llama2-7b) with a certain degree of mathematical problem-solving ability, [^2] Otherwise, the weak model can hardly provide useful supervision. denoted as $m$. This model acts analogously to human supervisors with limited expertise in the era of superintelligence. Besides, we only have a set of questions $\mathcal{Q}=\{q_{i}\}$ without ground truth answers and the goal is to improve the reasoning capability of a strong model $\mathcal{M}$ (e.g., Llama2-70b). To implement this, following Burns et al. (2023), we randomly divide the original training set into two equal parts, $\mathcal{D}_{\text{gold},1}$ and $\mathcal{D}_{\text{gold},2}$. The weak model is initially fine-tuned using $\mathcal{D}_{\text{gold},1}$ where the gold solutions are available, resulting in a weak model with some problem-solving capability, i.e. $m$. In contrast, the strong model can only access the questions from $\mathcal{D}_{\text{gold},2}$, without reasoning chains or final answers, i.e., $\mathcal{Q}$.
@@ 3.1.3 Weak-ICL Fine-Tuning
Given that models can mimic weak errors through supervised fine-tuning (Charikar et al., 2024; Lang et al., 2024), we propose refining $\mathcal{D}_{\text{weak}}$ before use, instead of using all data blindly. Additionally, we seek to harness the innate abilities of the strong model activated via in-context learning. Building on these two ideas, we introduce weak-icl fine-tuning, employing both weak data $\mathcal{D}_{\text{weak}}$ and “icl data” $\mathcal{D}_{\text{icl}}=\{q_{i},c_{\text{icl},i},a_{\text{icl},i}\}$, where $q_{i}\in\mathcal{Q}$, $c_{\text{icl},i}$ and $a_{\text{icl},i}$ are generated by $\mathcal{M}$ with few-shot demonstrations, [^3] Experiments in §4.3 show that despite ICL being affected by demonstration selection, our method can achieves further improvements accordingly beyond ICL. as higher-quality supervision signals.
@@ 4.3 Results of Stage I
The main results of Stage I on both GSM8K and MATH datasets are depicted in Fig. 4. Notably, in the MATH experiments, we randomly sample additional data that is not chosen based on the final answer consistency, due to the small amount available. Please refer to §A.4.1 for details. According to Fig. 4, we have the following observations.
@@ 4.6 Experiments Closer to Future Scenarios
Table 6: Results on OlympicArena using Llama3 family. The best result is in bold, and the best result of supervised fine-tuning in underlined.
@@
We only consider English questions in OlympicArena, excluding the CODE (Code Generation) and OT (Others) problem types that require case-based or expert evaluation. This results in 6,020 training data without solutions and final answers, and 313 test data with final answers to assess the performance of different methods. We use Llama3-8b-instruct (without initial fine-tuning on a subset of training data) as the weak model and Llama3-70b as the strong model to be improved. The hyperparameters are consistent with those used for GSM8K. This configuration more closely resembles future real-world weak-to-strong scenarios.
@@ 5.1 LLM Training
LLMs can enhance their ability to solve tasks and better align with human instructions through a supervised fine-tuning (SFT) phase (Zhang et al., 2023; Dong et al., 2023a; Lv et al., 2023b, a). This phase heavily relies on the quality of training data, as previous studies (Zhou et al., 2023a; Wang et al., 2023b) demonstrate that higher data quality translates to substantial gains in model performance. In this paper, we investigate the potential of learning from weak supervisions.
@@ 5.2 Mathematical Reasoning
The exploration of mathematical reasoning within LLMs has been a focal point for evaluating their cognitive capabilities akin to human reasoning (Qiao et al., 2023; Lu et al., 2023). Researchers have developed various methods to enhance the mathematical reasoning capabilities of LLMs after pre-training, which can be broadly classified into two categories: (1) Prompting: Some works (Kojima et al., 2022; Wei et al., 2022; Zhou et al., 2023b; Liu et al., 2023) aims to elicit the intrinsic reasoning abilities of LLMs by specific prompting engineering, without updating the model parameters; (2) Fine-tuning: Another line of studies focuses on generating a more extensive and higher-quality collection of question-answer pairs (Yu et al., 2023; Wang et al., 2023c, a). Through supervised fine-tuning and preference optimization (Luo et al., 2023; Azerbayev et al., 2023; Mitra et al., 2024; Xu et al., 2024), the models can achieve significant improvements in their mathematical problem-solving capabilities.
@@ 6 Conclusion
In this paper, we explore the efficacy of weak-to-strong framework in complex reasoning tasks. We introduce a new method that elicits strong capabilities using weak supervisions, without relying on annotations from humans or more advanced models. This method focuses on the strong model’s ability to autonomously refine its training data, even if it has not learned the task before. By iteratively expanding its learning scope, the strong model continuously broadens its reasoning skills. This self-directed data curation is crucial for scaling up the enhancement of AI reasoning capabilities, making the model more independent and effective in its developmental trajectory. Through this work, we seek to illuminate new pathways for AI development, emphasizing the critical role of innovative model supervision in advancing AGI and beyond.
@@ Acknowledgements
We sincerely thank Xuefeng Li, Haoyang Zou, and Ting Wu for their valuable insights during discussions, which greatly enhance the quality of this work.
@@ A.2 Training Details
For supervised fine-tuning in Stage I, we adopt LoRA to fine-tune the strong model $\mathcal{M}$ with a learning rate of $1\times 10^{-4}$ and search for weight decay in the set $\{0,0.01\}$. We run 2 epochs on GSM8K and 3 epochs on MATH, with a batch size of 8. In Stage II, we employ two preference optimization methods. For DPO, we train the enhanced strong model $\mathcal{M}_{\text{plus}}$ with a learning rate of $1\times 10^{-5}$ and run 1 epoch. For ORPO, we search for $\beta$ in the set $\{0.1,0.5,1.0\}$ with a learning rate of $3\times 10^{-5}$ and run 1 epoch. All experiments are conducted using A100 GPUs.
When constructing contrastive samples in Stage II, we sample $n=10$ responses at $\text{temperature}=1.0$, and use a confidence threshold of $\tau=0.6$. Normally, we evaluate using greedy decoding. For calculating pass@k, we set $k=10$ at $\text{temperature}=1.0$.
@@ A.4.4 Effect of SFT Data
Question:
{question}
Student Solution:
{solution}
Your task involves three parts:
Here’s the format I want:
Final Judgement: [Insert only **correct** or **wrong** here]
Table 13: Prompt used to evaluate process-level accuracy.

### v2

@@ Abstract
When large language models (LLMs) exceed human-level capabilities, it becomes increasingly challenging to provide full-scale and accurate supervision for these models. Weak-to-strong learning, which leverages a less capable model to unlock the latent abilities of a stronger model, proves valuable in this context. Yet, the efficacy of this approach for complex reasoning tasks is still untested. Furthermore, tackling reasoning tasks under the weak-to-strong setting currently lacks efficient methods to avoid blindly imitating the weak supervisor including its errors. In this paper, we introduce a progressive learning framework that enables the strong model to autonomously refine its training data, without requiring input from either a more advanced model or human-annotated data. This framework begins with supervised fine-tuning on a selective small but high-quality dataset, followed by preference optimization on contrastive samples identified by the strong model itself. Extensive experiments on the GSM8K and MATH datasets demonstrate that our method significantly enhances the reasoning capabilities of Llama2-70b using three separate weak models. This method is further validated in a forward-looking experimental setup, where Llama3-8b-instruct effectively supervises Llama3-70b on the highly challenging OlympicArena dataset. This work paves the way for a more scalable and sophisticated strategy to enhance AI reasoning powers. All relevant code and resources are available in https://github.com/GAIR-NLP/weak-to-strong-reasoning.
@@ 1 Introduction
To achieve the above goal, we introduce a progressive refinement learning framework, guided by the principle that a model can enhance its capabilities more effectively by initially focusing on smaller, more reliable subsets of data, and then iteratively expanding its learning scope, as illustrated in Fig. 2. In the first stage, we hypothesize that it is more advantageous to utilize smaller quantities of data that are likely to be more accurate. We achieve this by combining weak data, generated by the less capable model, with data self-generated by the more advanced model through in-context learning. This blend is then used to selectively curate datasets for subsequent supervised fine-tuning. In the second stage, upon having developed a strong model with improved reasoning capabilities, we utilize its ability to construct contrastive samples for preference optimization (Rafailov et al., 2023; Hong et al., 2024) and enable the model to learn effectively from the errors of the weaker model.
@@
To more accurately approximate future scenarios, we conduct additional experiments on OlympicArena (Huang et al., 2024), an extremely challenging dataset with no definitive ground truth answers. Llama3-8b-instruct (AI@Meta, 2024), despite its smaller size, has been aligned and proven to effectively supervise the larger Llama3-70b, whose potential has not yet been fully realized. Moreover, our proposed two-stage training approach outperforms full weak fine-tuning by 3.19 points.
@@ G.T. Answer Stronger Model
Gaining insights from semi-supervised learning within the domain of traditional machine learning, another type of LLM learning depends not on extensive labeling but instead on a small, high-quality seed dataset. Tong et al. (2024) have demonstrated improvement by learning differences between self-generated responses and expert-annotated responses. We also include the trending research topic of easy-to-hard generalization (Hase et al., 2024; Sun et al., 2024) in this category, where models are trained to tackle complex tasks by learning from human annotations on easier tasks. This series of research inevitably requires access to a small yet high-quality set of standard answers.
@@ 2.2 Weak-to-Strong Reasoning Setup
In this paper, we address reasoning tasks in the weak-to-strong setting, as illustrated in Tab. 2. First, we examine mathematical reasoning tasks, such as those in GSM8k and MATH. These tasks require each step of the reasoning process to demonstrate fundamental mathematical problem-solving skills, including problem comprehension and algebraic operations, and build upon the previous steps. This imposes higher demands on the model’s learning and generalization capabilities. Unlike classification tasks, where models can rely on superficial pattern extrapolation or recognition, reasoning tasks offer minimal benefit from guessing. Then, we use a weak model (e.g., Llama2-7b) with a certain degree of mathematical problem-solving ability, [^2] Otherwise, the weak model can hardly provide useful supervision. denoted as $m$. This model acts analogously to human supervisors with limited expertise in the era of superintelligence. Besides, we only have a set of questions $\mathcal{Q}=\{q_{i}\}$ without ground truth answers and the goal is to improve the reasoning capability of a strong model $\mathcal{M}$ (e.g., Llama2-70b). To implement this, following Burns et al. (2023), we randomly divide the original training set into two equal parts, $\mathcal{D}_{\text{gold},1}$ and $\mathcal{D}_{\text{gold},2}$. The weak model is initially fine-tuned using $\mathcal{D}_{\text{gold},1}$ where the gold solutions are available, resulting in a weak model with some problem-solving capability, i.e. $m$. In contrast, the strong model can only access the questions from $\mathcal{D}_{\text{gold},2}$, without reasoning chains or final answers, i.e., $\mathcal{Q}$.
@@ 3.1.3 Weak-ICL Fine-Tuning
Given that models can mimic weak errors through supervised fine-tuning (Charikar et al., 2024; Lang et al., 2024), we propose refining $\mathcal{D}_{\text{weak}}$ before use, instead of using all data blindly. Additionally, we seek to harness the innate abilities of the strong model activated via in-context learning. Building on these two ideas, we introduce weak-icl fine-tuning, employing both weak data $\mathcal{D}_{\text{weak}}$ and “icl data” $\mathcal{D}_{\text{icl}}=\{q_{i},c_{\text{icl},i},a_{\text{icl},i}\}$, where $q_{i}\in\mathcal{Q}$, $c_{\text{icl},i}$ and $a_{\text{icl},i}$ are generated by $\mathcal{M}$ with few-shot demonstrations, [^3] Experiments in §4.3 show that despite ICL being affected by demonstration selection, our method can achieve further improvements accordingly beyond ICL. as higher-quality supervision signals.
@@ 4.3 Results of Stage I
The main results of Stage I on both GSM8K and MATH datasets are depicted in Fig. 4.444We do not incorporate the zero-shot performance of the strong model in the main results as it is significantly lower than that of weak ICL. See §A.4.5 for further details. Notably, in the MATH experiments, we randomly sample additional data that is not chosen based on the final answer consistency, due to the small amount available. Please refer to §A.4.1 for details. According to Fig. 4, we have the following observations.
@@ 4.6 Experiments Closer to Future Scenarios
Table 6: Results on OlympicArena using Llama3 family. The best result is in bold, and the best result of supervised fine-tuning is underlined.
@@
We only consider English questions in OlympicArena, excluding the CODE (Code Generation) and OT (Others) problem types that require case-based or expert evaluation. This results in 6,020 training data without solutions and final answers, and 313 test data with final answers to assess the performance of different methods. We use Llama3-8b-instruct (without initial fine-tuning on a subset of training data) as the weak model and Llama3-70b as the strong model to be improved. This configuration more closely resembles future real-world weak-to-strong scenarios.
@@ 5.1 LLM Training
LLMs can enhance their ability to solve tasks and better align with human instructions through a supervised fine-tuning (SFT) phase (Zhang et al., 2023; Dong et al., 2023a; Lv et al., 2023b, a). This phase heavily relies on the quality of training data, as previous studies (Zhou et al., 2023a; Wang et al., 2023b) demonstrate that higher data quality translates to substantial gains in model performance. In this paper, we investigate the potential of learning from weak supervision.
@@ 5.2 Mathematical Reasoning
The exploration of mathematical reasoning within LLMs has been a focal point for evaluating their cognitive capabilities akin to human reasoning (Qiao et al., 2023; Lu et al., 2023). Researchers have developed various methods to enhance the mathematical reasoning capabilities of LLMs after pre-training, which can be broadly classified into two categories: (1) Prompting: Some works (Kojima et al., 2022; Wei et al., 2022; Zhou et al., 2023b; Liu et al., 2023) aim to elicit the intrinsic reasoning abilities of LLMs by specific prompting engineering, without updating the model parameters; (2) Fine-tuning: Another line of studies focuses on generating a more extensive and higher-quality collection of question-answer pairs (Yu et al., 2023; Wang et al., 2023c, a). Through supervised fine-tuning and preference optimization (Luo et al., 2023; Azerbayev et al., 2023; Mitra et al., 2024; Xu et al., 2024), the models can achieve significant improvements in their mathematical problem-solving capabilities.
@@ 6 Conclusion
In this paper, we explore the efficacy of weak-to-strong framework in complex reasoning tasks. We introduce a new method that elicits strong capabilities using weak supervision, without relying on annotations from humans or more advanced models. This method focuses on the strong model’s ability to autonomously refine its training data, even if it has not learned the task before. By iteratively expanding its learning scope, the strong model continuously broadens its reasoning skills. This self-directed data curation is crucial for scaling up the enhancement of AI reasoning capabilities, making the model more independent and effective in its developmental trajectory. Through this work, we seek to illuminate new pathways for AI development, emphasizing the critical role of innovative model supervision in advancing AGI and beyond.
@@ Acknowledgements
We sincerely thank Xuefeng Li, Haoyang Zou, and Ting Wu for their valuable insights during discussions, which greatly enhanced the quality of this work. This work was supported by Shanghai Artificial Intelligence Laboratory, SJTU SEIEE - ByteDance Large Language Model Joint Laboratory.
@@ A.2 Training Details
For supervised fine-tuning in Stage I, we adopt LoRA to fine-tune the strong model $\mathcal{M}$ with a learning rate of $1\times 10^{-4}$ and search for weight decay in the set $\{0,0.01\}$. We run 2 epochs on GSM8K and 3 epochs on MATH, with a batch size of 8. In Stage II, we employ two preference optimization methods. For DPO, we train the enhanced strong model $\mathcal{M}_{\text{plus}}$ with a learning rate of $1\times 10^{-5}$ and run 1 epoch. For ORPO, we search for $\beta$ in the set $\{0.1,0.5,1.0\}$ with a learning rate of $3\times 10^{-5}$ and run 1 epoch. For experiments on OlympicArena using Llama3 family, the hyperparameters are consistent with those used for GSM8K. All experiments are conducted using A100 GPUs. Moreover, when constructing contrastive samples in Stage II, we sample $n=10$ responses at $\text{temperature}=1.0$, and use a confidence threshold of $\tau=0.6$. Normally, we evaluate using greedy decoding. For calculating pass@k, we set $k=10$ and $\text{temperature}=1.0$.
@@ A.3.2 Training Accuracy of Stage I
Question:
{question}
Student Solution:
{solution}
Your task involves three parts:
Here’s the format I want:
Final Judgement: [Insert only **correct** or **wrong** here]
Table 9: Prompt used to evaluate process-level accuracy.
@@ A.4.4 Effect of SFT Data
A.4.5 Zero-Shot Results
Test Accuracy
GSM8K
Llama2-70b wo CoT 12.36
Llama2-70b w/ CoT 18.35
MATH
Llama2-70b wo CoT 6.40
Llama2-70b w/ CoT 7.20
Table 14: Zero-Shot Results of Llama2-70b on GSM8K and MATH.
[…]